# in World Models

### v1

@@ 1 Introduction
We present MIND, the first closed-loop revisited open-domain benchmark for evaluating memory consistency and action control from both first-person and third-person perspectives across diverse scenarios. MIND focuses on two key abilities of world models: 1) Memory consistency refers to the ability of model to maintain coherent spatial layouts, object identities, and scene attributes over long temporal contexts, ensuring that generated frames remain consistent with past observations. 2) Action control measures how accurately the model executes given control inputs and generalizes these dynamics to new motion ranges or unseen action spaces, reflecting its capacity for precise and adaptable interaction within dynamic environments. Furthermore, the provided videos include frame-level aligned actions, character and camera positions, and image labels, collected from multiple volunteers to capture diverse human behaviors. The dataset contains $250$ high-quality 10801080p, $24$ FPS, frame-level action-aligned videos spanning eight major scene categories, enabling comprehensive evaluation of world models.
@@
  • Evaluation for Cross-Action Space Generalization. We design various action spaces, including different character movement speeds and camera rotation angles, to evaluate the action generalization capability across different action spaces under shared scenes.
@@ 2.1 Video Generation
Recent advances in video generation models such as SVD [1], Hunyuanvideo [20], Cogvideox [45], Wan [40] and Sora 2 [31] have significantly enhanced video realism, temporal coherence, and controllability, extending generation toward long-horizon, physically plausible scenes. While benchmarks such as VBench [18], VBench-2.0 [54] and EvalCrafter [27] have introduced fine-grained evaluation dimensions (e.g., human fidelity, physics, commonsense).

### v2

@@ 1 Introduction
We present MIND, the first closed-loop revisited open-domain benchmark for evaluating memory consistency and action control from both first-person and third-person perspectives across diverse scenarios. MIND focuses on two key abilities of world models: 1) Memory consistency refers to the ability of model to maintain coherent spatial layouts, object identities, and scene attributes over long temporal contexts, ensuring that generated frames remain consistent with past observations. 2) Action control measures how accurately the model executes given control inputs and generalizes these dynamics to new motion ranges or unseen action spaces, reflecting its capacity for precise and adaptable interaction within dynamic environments. Furthermore, the provided videos include frame-level aligned actions, character and camera positions, and image labels, collected from multiple volunteers to capture diverse human behaviors. The dataset contains $250$ high-quality 10801080p / $24$ FPS, frame-level action-aligned videos spanning eight major scene categories, enabling comprehensive evaluation of world models.
@@
  • Evaluation for Cross-Action Space Generalization. We design various action spaces, including different character movement speeds and camera rotation angles, to evaluate the action space generalization capability.
@@ 2.1 Video Generation
Recently, video generation models represented by SVD [1], Hunyuanvideo [20], Wan [39] and Sora 2 [30] have achieved remarkable breakthroughs. While significantly enhancing the visual realism, temporal consistency and controllability of generated videos, these models have extended their generation capabilities to long-sequence and physically plausible scenarios. Benchmarks including VBench [18] and VBench-2.0 [52] have established dedicated evaluation systems for video generation models, covering key dimensions such as human fidelity, physical plausibility and commonsense consistency.